# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products with applications to Kohn-Sham density functional theory

### v4

@@ 4.1 Real Symmetric Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at the origin (Gamma-points) of the Brillouin zone for the systems described in table 1. This ensures that the resulting discretised equation is a real symmetric eigenvalue problem.
@@
From fig. 1, we observe that for all the benchmark systems summarised in table 1 and various Chebyshev polynomial degrees, the residual tolerance that can be achieved by the R-ChFSI method is orders of magnitude lower than what can be achieved using the ChFSI method when employing FP64 arithmetic and the diagonal approximation for the inverse overlap matrix in the subspace filtering step. This observation is consistent with the analysis done in section 3. As previously discussed, the R-ChFSI algorithm enables us to construct the filtered subspace using lower precision arithmetic. We report the residual norms ($\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$) achieved by the FP32, TF32 and TF32B variants of R-ChFSI in fig. 3. We note that the residual norms obtained using the FP32, TF32 and TF32B variants are comparable to those obtained using the FP64 variant for Chebyshev polynomial degrees of $p=20,40,60,80$ but not for $p=100,120$. We believe this is because the condition described in eq. 5 might not hold for higher values of $p$.
@@ 4.2 Complex Hermitian Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at a non-zero $k$-point  [dasDFTFE10Massively2022] in the Brillouin zone for the systems described in table 1. The resulting discretized equation is now a complex Hermitian eigenvalue problem.
@@
We also note that our proposed R-ChFSI algorithm allows us to construct the filtered subspace using lower precision for the complex Hermitian eigenproblem, and we report the residuals achieved by the FP32, TF32 and TF32B variants of R-ChFSI in fig. 6. We note that the residual norms obtained using the FP32, TF32 and TF32B variants are comparable to those obtained using the FP64 variant for Chebyshev polynomial degrees of $p=20,40,60,80$.
@@
We also report the speedups achieved for the filtering step for Chebyshev polynomial degrees of $p=40,60,80$ by the FP32, TF32 and TF32B variants of R-ChFSI over the FP64 variant of R-ChFSI in fig. 7 on Intel Data Center GPU Max Series accelerators deployed in the Aurora supercomputing system. We achieve speedups of up to 2.3x for the TF32 variant and 2.7x for the TF32B variant of the R-ChFSI method over the FP64 variant of the R-ChFSI method. We further report the speedups achieved for the total eigensolve (with the convergence criteria $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,<10^{-8}$) for Chebyshev polynomial degrees of $p=40,60,80$ by the FP32, TF32 and TF32B variants of R-ChFSI over the FP64 variant of R-ChFSI in fig. 8 on Intel Data Center GPU Max Series accelerators deployed in the Aurora supercomputing system. For the full eigensolve, we achieve speedups of up to 1.9x for the TF32 and 2.1x for the TF32B variant of the R-ChFSI method over the FP64 variant of the R-ChFSI method.
@@ 5 Conclusion
In this work, we have proposed a residual-based Chebyshev filtered subspace iteration (R-ChFSI) method designed for large-scale sparse Hermitian eigenvalue problems, with a particular emphasis on robustness to inexact matrix–vector products. Through both mathematical analysis and extensive numerical experiments, we have demonstrated that the proposed method outperforms the standard Chebyshev filtered subspace iteration (ChFSI) approach significantly. Our theoretical justification demonstrates how reformulating the Chebyshev recurrence to operate on residuals, rather than directly on eigenvector estimates, effectively suppresses numerical errors introduced by inexact operator applications. This leads to a more reliable construction of the filtered subspace when compared with the traditional ChFSI recurrence relation.
An important contribution of this work lies in demonstrating the effectiveness of R-ChFSI for generalized Hermitian eigenvalue problems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$. In this setting, the proposed method is particularly impactful due to its ability to incorporate approximate inverses of B when constructing the subspace rich in the desired eigenspace using the Chebyshev polynomial recurrence relation. By replacing the exact computation of matrix inverses with computationally inexpensive approximate inverses (e.g., diagonal or block-diagonal approximations of B), the method not only reduces the overall computational burden but also maintains convergence robustness. This results in substantial performance gains, as the proposed approach achieves residual tolerances that are orders of magnitude lower than those obtainable with traditional ChFSI when solving the generalized problems. This highlights the method’s robustness and its suitability for large-scale discretised sparse generalized eigenproblems where exact factorisations of B are prohibitively expensive.
@@
The proposed R-ChFSI methodology has the potential to significantly impact a wide range of computational physics applications. These include ab initio modelling of materials, structural mechanics, fluid dynamics, and multiphysics problems, all of which rely on the efficient solution of large, sparse Hermitian eigenvalue problems arising from discretisations of partial differential equations using finite-element, finite-difference, wavelets or finite-volume formulations. By enabling the effective use of low-precision arithmetic without sacrificing accuracy, the method provides a compelling alternative for large-scale simulations that require substantial computational resources. Moreover, its robustness in solving large generalized nonlinear eigenvalue problems opens new opportunities for applying R-ChFSI to domains such as quantum modeling of materials, and other areas that demand scalable and reliable sparse eigensolvers.
@@ 7 Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used Paperpal in order to proofread. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ A Useful Lemmas
Let the matrices $\textbf{E}\in\mathbb{C}^{m\times n}$ and $\textbf{E}_{\perp}\in\mathbb{C}^{m\times(m-n)}$ be given by $\textbf{E}=\Bigl[\begin{smallmatrix}\textbf{I}_{n}\\ 0\end{smallmatrix}\Bigr]$ and $\textbf{E}_{\perp}=\Bigl[\begin{smallmatrix}0\\ \textbf{I}_{m-n}\end{smallmatrix}\Bigl]$ where $\textbf{I}_{n}$ and $\textbf{I}_{m-n}$ are the $n\times n$ and $(m-n)\times(m-n)$ identity matrices respectively. Let $\mathcal{E}=\mathcal{R}(\textbf{E})$ and $\mathcal{E}_{\perp}=\mathcal{R}(\textbf{E}_{\perp})$ where $\mathcal{R}(.)$ denotes range of a matrix. Then for any matrix $\textbf{X}\in\mathbb{C}^{m\times n}$ with $\mathcal{X}=\mathcal{R}(\textbf{X})$ satisfying $\mathcal{E}_{\perp}\cap\mathcal{X}=\left\{0\right\}$ or equivalently $\text{rank}\left(\textbf{E}^{\dagger}\textbf{X}\right)=n$, we have
$\tan{\angle(\mathcal{X},\mathcal{E})}=\lVert\textbf{E}_{\perp}^{\dagger}\textbf{X}(\textbf{E}^{\dagger}\textbf{X})^{-1}\rVert\,$
Proof.
For proof of this statement, we refer the reader to [zhuAnglesSubspacesTheir2013, watkinsMatrixEigenvalueProblem2007]. ∎
Lemma A.2.
@@
This is a straightforward application of the results derived in [highamAccuracyStabilityNumerical2002, section 3.5]. We note that $\gamma_{m}$ is a constant that depends on $m$, and the machine precision of the floating point arithmetic (denoted as $\varepsilon_{M}$). For the case of dense real matrices, we have $\gamma_{m}=m^{2}\varepsilon_{M}/(1-m\varepsilon_{M})$ and for dense complex matrices, it needs to be appropriately modified as described in [highamAccuracyStabilityNumerical2002, section 3.6]. We also note that tigher bounds can be obtained from probabilistic error analysis [ipsenProbabilisticErrorAnalysis2020, highamSharperProbabilisticBackward2020]. Further, we also note that in the case where H is sparse, significantly tighter bounds independent of $m$ can be achieved depending on the specific implementations of the sparse matrix-dense matrix multiplication routines. ∎
@@
$\textbf{d}_{k}^{(i)}$ $=\begin{bmatrix}\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\\ \lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,\end{bmatrix}$ $\textbf{F}_{k}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,+\lvert b_{k}\rvert\,+\lvert a_{k}\rvert\,h_{0}&\lvert c_{k}\rvert\,\\ 1&0\end{bmatrix}$ $\textbf{E}_{k}^{(i)}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,h_{1}^{(i)}\\ 0,\end{bmatrix}$
@@
Using submultiplicativity of the spectral norm we can now write
$\lVert\textbf{R}_{k}^{(i)}\rVert\,$ $\leq\sum_{j=0}^{k}\sum_{r=0}^{j-1}\lVert\alpha_{j}\textbf{H}^{j-r-1}\rVert\,\lVert\textbf{H}{\textbf{X}^{(i)}}-{\textbf{X}^{(i)}}{\boldsymbol{\Lambda}^{(i)}}\rVert\,\lVert{\boldsymbol{\Lambda}^{(i)}}^{r}\rVert\,=f_{k}\lVert\textbf{R}^{(i)}\rVert\,$
where we have defined $f_{k}=\sum_{j=0}^{k}\sum_{r=0}^{j-1}\lVert\alpha_{j}\textbf{H}^{j-r-1}\rVert\,\lVert{\boldsymbol{\Lambda}^{(i)}}^{r}\rVert\,$. ∎
@@
Using the fact that $\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}={\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{1}\rVert\,=\lVert\hat{\textbf{U}}_{2}^{\dagger}\textbf{X}^{(i)}\rVert\,}$ and the inequalities $\lVert\hat{\textbf{U}}_{1}^{\dagger}\textbf{X}^{(i)}\rVert\,\leq 1$, $\lVert(\textbf{I}-{\textbf{X}^{(i)}}^{\dagger}\textbf{X}^{(i)})\hat{\textbf{U}}_{2}\rVert\,\leq 1$ we can write
@@
with the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=0$ and $\boldsymbol{\Delta}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{X}^{(i)}$, obtained from the initial conditions of eq. 2 and eq. 7. While $\textbf{e}_{k}^{(i)}=\textbf{D}^{-1}\otimes\textbf{A}\otimes\underline{\textbf{Y}}_{k}^{(i)}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}$. From lemma A.2 we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,=\lVert{\textbf{D}^{-1}\otimes\textbf{A}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}\rVert\,\leq(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,$. We further note that
@@
Thus we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$. Using lemma A.3 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=0$ and $\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,=\lVert\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{X}^{(i)}\rVert\,$, $h_{0}=(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,$ and $h_{1}^{(i)}=(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k}$ with $\eta_{k}=2\lVert\textbf{D}^{(-1)}\rVert\,\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,g_{k-1}^{1}$ and $\tilde{\eta}_{k}=\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,g_{k-1}^{1}+g_{k-1}^{0}\lvert\frac{\sigma_{1}}{e}\rvert\,$.
@@
We further note that using lemma A.4 we can write

### v5

@@ 4.2 Real Symmetric Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at the origin (Gamma-points) of the Brillouin zone for the systems described in table 1. This ensures that the resulting discretized equation is a real symmetric eigenvalue problem.
@@
From fig. 4, we observe that for all the benchmark systems summarized in table 1 and various Chebyshev polynomial degrees, the residual tolerance that can be achieved by the R-ChFSI method is orders of magnitude lower than what can be achieved using the ChFSI method when employing FP64 arithmetic and the diagonal approximation for the inverse overlap matrix in the subspace filtering step. This observation is consistent with the analysis done in section 3. As previously discussed, the R-ChFSI algorithm enables us to construct the filtered subspace using lower precision arithmetic. We report the residual norms ($\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$) achieved by the FP32, TF32 and TF32B variants of R-ChFSI in fig. 6. We note that the residual norms obtained using the FP32, TF32 and TF32B variants are comparable to those obtained using the FP64 variant for Chebyshev polynomial degrees of $p=20,40,60,80$ but not for $p=100,120$. As demonstrated by our controlled experiments in section 4.1.4, this can be attributed to the convergence condition of eq. 5 being violated at higher polynomial degrees, where the accumulated filtering error $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ exceeds the left-hand side $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$.
@@ 4.3 Complex Hermitian Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at a non-zero $k$-point  [8] in the Brillouin zone for the systems described in table 1. The resulting discretized equation is now a complex Hermitian eigenvalue problem.
@@
We also note that our proposed R-ChFSI algorithm allows us to construct the filtered subspace using lower precision for the complex Hermitian eigenproblem, and we report the residuals achieved by the FP32, TF32 and TF32B variants of R-ChFSI in fig. 9. We note that the residual norms obtained using the FP32, TF32 and TF32B variants are comparable to those obtained using the FP64 variant for Chebyshev polynomial degrees of $p=20,40,60,80$. We also report the speedups achieved for the filtering step for Chebyshev polynomial degrees of $p=40,60,80$ by the FP32, TF32 and TF32B variants of R-ChFSI over the FP64 variant of R-ChFSI in fig. 10 on Intel Data Center GPU Max Series accelerators deployed in the Aurora supercomputing system. We achieve speedups of up to 2.3x for the TF32 variant and 2.7x for the TF32B variant of the R-ChFSI method over the FP64 variant of the R-ChFSI method. We further report the speedups achieved for the total eigensolve (with the convergence criteria $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,<10^{-8}$) for Chebyshev polynomial degrees of $p=40,60,80$ by the FP32, TF32 and TF32B variants of R-ChFSI over the FP64 variant of R-ChFSI in fig. 11 on Intel Data Center GPU Max Series accelerators deployed in the Aurora supercomputing system. For the full eigensolve, we achieve speedups of up to 1.9x for the TF32 and 2.1x for the TF32B variant of the R-ChFSI method over the FP64 variant of the R-ChFSI method.
@@
4.3.1 Summary: complex Hermitian vs. real symmetric
Comparing the complex Hermitian results above with the real symmetric results in the preceding subsection, three observations stand out. First, the accuracy benefit of R-ChFSI over ChFSI is equally pronounced in both cases: the residual-based reformulation reduces the stagnation floor by several orders of magnitude regardless of whether the matrices are real or complex, confirming that the theoretical guarantees of section 3 hold without modification for the complex Hermitian setting. Second, the lower-precision variants (FP32, TF32, TF32B) remain as robust for complex matrices as for real ones, converging to the same residual tolerances at polynomial degrees $p\leq 80$. Third, the performance speedups are uniformly larger in the complex case (up to $2.3{\times}$ for TF32 and $2.7{\times}$ for TF32B in the filtering step, versus $1.8{\times}$ and $2.6{\times}$ in the real case). This is expected: complex arithmetic doubles the number of floating-point operations and the volume of data per matrix element, so halving the precision yields a proportionally greater reduction in both memory traffic and computation. The full-eigensolve speedups ($1.9{\times}$$2.1{\times}$ for complex vs. $1.6{\times}$$2.0{\times}$ for real) follow the same trend, with the slightly smaller gains reflecting the fixed FP64 cost of the Rayleigh–Ritz projection. Overall, these results demonstrate that R-ChFSI with low-precision arithmetic is effective across the full range of DFT calculations—both $\Gamma$-point (real symmetric) and $k$-point (complex Hermitian).
@@ 5 Conclusion
In this work, we have proposed a residual-based Chebyshev filtered subspace iteration (R-ChFSI) method designed for large-scale Hermitian eigenvalue problems, with a particular emphasis on robustness to inexact matrix–vector products. Through both mathematical analysis and extensive numerical experiments, we have demonstrated that the proposed method outperforms the standard Chebyshev filtered subspace iteration (ChFSI) approach significantly. Our theoretical justification demonstrates how reformulating the Chebyshev recurrence to operate on residuals, rather than directly on eigenvector estimates, effectively suppresses numerical errors introduced by inexact operator applications. This leads to a more reliable construction of the filtered subspace when compared with the traditional ChFSI recurrence relation.
An important contribution of this work lies in demonstrating the effectiveness of R-ChFSI for generalized Hermitian eigenvalue problems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$. In this setting, the proposed method is particularly impactful due to its ability to incorporate approximate inverses of B when constructing the subspace rich in the desired eigenspace using the Chebyshev polynomial recurrence relation. By replacing the exact computation of matrix inverses with computationally inexpensive approximate inverses (e.g., diagonal or block-diagonal approximations of B), the method not only reduces the overall computational burden but also maintains convergence robustness. This results in substantial performance gains, as the proposed approach achieves residual tolerances that are orders of magnitude lower than those obtainable with traditional ChFSI when solving the generalized problems. This highlights the method’s robustness and its suitability for large-scale discretized generalized eigenproblems where exact factorizations of B are prohibitively expensive.
@@
The proposed R-ChFSI methodology has the potential to significantly impact a wide range of computational physics applications. These include ab initio modeling of materials, structural mechanics, fluid dynamics, and multiphysics problems, all of which rely on the efficient solution of large Hermitian eigenvalue problems arising from discretizations of partial differential equations using finite-element, finite-difference, wavelets or finite-volume formulations. By enabling the effective use of low-precision arithmetic without sacrificing accuracy, the method provides a compelling alternative for large-scale simulations that require substantial computational resources. Moreover, its robustness in solving large generalized nonlinear eigenvalue problems opens new opportunities for applying R-ChFSI to domains such as quantum modeling of materials, and other areas that demand scalable and reliable eigensolvers.
@@ 7 Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used Paperpal in order to proofread and Claude (Anthropic, via github copilot) in order to assist with coding for dense eigenproblem testing. After using these tools/services, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
8 Data availability
The code and data associated with this article can be found online at https://github.com/matrixlabiisc/R-ChFSI-Manuscript-Data.
@@ A Useful Lemmas
This is a straightforward application of the results derived in [16, section 3.5]. We note that $\gamma_{m}$ is a constant that depends on $m$, and the machine precision of the floating point arithmetic (denoted as $\varepsilon_{M}$). For the case of dense real matrices, we have $\gamma_{m}=m^{2}\varepsilon_{M}/(1-m\varepsilon_{M})$ and for dense complex matrices, it needs to be appropriately modified as described in [16, section 3.6]. We also note that tighter bounds can be obtained from probabilistic error analysis [18, 15]. Further, we also note that in the case where H is sparse, significantly tighter bounds independent of $m$ can be achieved depending on the specific implementations of the sparse matrix-dense matrix multiplication routines. ∎
@@
$\textbf{d}_{k}^{(i)}$ $=\begin{bmatrix}\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\\ \lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,\end{bmatrix}$ $\textbf{F}_{k}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,+\lvert b_{k}\rvert\,+\lvert a_{k}\rvert\,h_{0}&\lvert c_{k}\rvert\,\\ 1&0\end{bmatrix}$ $\textbf{E}_{k}^{(i)}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,h_{1}^{(i)}\\ 0\end{bmatrix}$
@@
Since $\boldsymbol{\Lambda}^{(i)}$ comprises Ritz values of the Hermitian matrix H, each diagonal entry of $\boldsymbol{\Lambda}^{(i)}$ lies in the interval $\lambda_{1},\lambda_{m}]$ and hence $\lVert\boldsymbol{\Lambda}^{(i)}\rVert\,=\max_{j}|\epsilon_{j}^{(i)}|\leq\max\{|\lambda_{1}|,|\lambda_{m}|\}\leq\lVert\textbf{H}\rVert\,$. Using submultiplicativity of the spectral norm we can now write
$\lVert\textbf{R}_{k}^{(i)}\rVert\,$ $\leq\sum_{j=1}^{k}\sum_{r=0}^{j-1}|\alpha_{j}|\lVert\textbf{H}\rVert\,^{j-r-1}\lVert\textbf{H}{\textbf{X}^{(i)}}-{\textbf{X}^{(i)}}{\boldsymbol{\Lambda}^{(i)}}\rVert\,\lVert\boldsymbol{\Lambda}^{(i)}\rVert\,^{r}\leq\sum_{j=1}^{k}\sum_{r=0}^{j-1}|\alpha_{j}|\lVert\textbf{H}\rVert\,^{j-1}\lVert\textbf{R}^{(i)}\rVert\,=f_{k}\lVert\textbf{R}^{(i)}\rVert\,$
where we have defined $f_{k}=\sum_{j=1}^{k}j|\alpha_{j}|\lVert\textbf{H}\rVert\,^{j-1}$. ∎
@@
Using the fact that $\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}={\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{1}\rVert\,=\lVert\hat{\textbf{U}}_{2}^{\dagger}\textbf{X}^{(i)}\rVert\,}$ and the inequalities $\lVert\hat{\textbf{U}}_{1}^{\dagger}\textbf{X}^{(i)}\rVert\,\leq 1$, $\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{2}\rVert\,\leq 1$ we can write
@@
with the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=0$ and $\boldsymbol{\Delta}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{A}\textbf{X}^{(i)}$, obtained from the initial conditions of eq. 2 and eq. 7. While $\textbf{e}_{k}^{(i)}=\textbf{D}^{-1}\otimes\textbf{A}\otimes\underline{\textbf{Y}}_{k}^{(i)}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}$. From lemma A.1 we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,=\lVert{\textbf{D}^{-1}\otimes\textbf{A}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}\rVert\,\leq(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,$. We further note that
@@
Thus we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$. Further, since $\hat{\textbf{X}}^{(i)}=\textbf{B}^{\frac{1}{2}}\textbf{X}^{(i)}$ is orthonormal, we have $\lVert\textbf{X}^{(i)}\rVert\,\leq\lVert\textbf{B}^{-\frac{1}{2}}\rVert\,$. Using lemma A.2 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=0$ and $\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\leq\lvert\frac{\sigma_{1}}{e}\rvert\,\zeta\lVert\textbf{A}\rVert\,\lVert\textbf{B}^{-\frac{1}{2}}\rVert\,$, $h_{0}=(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,$ and $h_{1}^{(i)}=(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$ we obtain $\lVert\hat{\boldsymbol{\Delta}}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k}$ with $\eta_{k}=2\lVert\textbf{B}^{\frac{1}{2}}\rVert\,\lVert\textbf{D}^{-1}\rVert\,\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,g_{k-1}^{1}$ and $\tilde{\eta}_{k}=\lVert\textbf{B}^{\frac{1}{2}}\rVert\,(\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,g_{k-1}^{1}+g_{k-1}^{0}\lvert\frac{\sigma_{1}}{e}\rvert\,\lVert\textbf{A}\rVert\,\lVert\textbf{B}^{-\frac{1}{2}}\rVert\,)$.
@@
Further, since $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$ and $\lVert\boldsymbol{\Lambda}^{(i)}\rVert\,\leq\lVert\textbf{H}\rVert\,$, we have $\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,\leq\max_{|x|\leq\lVert\textbf{H}\rVert\,}|C_{k}(x)|=:M_{k}$. We further note that using lemma A.3 we can write